OpenStudy (anonymous):

which of these values for P and A will cause the function f(x)=Pa* to be an exponential growth functon? [* is an x variable]

OpenStudy (anonymous):

P=4; a=1 P=1/4; a=1/5 P=4; a=1/5 P=1/4; a=5

OpenStudy (rogue):

a must be greater than 1 for it to be an exponential growth function, so it is choice 4.